# Gradient Boosting Reinforcement Learning

### v1

@@ Abstract
Neural networks (NN) achieve remarkable results in various tasks, but lack key characteristics: interpretability, support for categorical features, and lightweight implementations suitable for edge devices. While ongoing efforts aim to address these challenges, Gradient Boosting Trees (GBT) inherently meet these requirements. As a result, GBTs have become the go-to method for supervised learning tasks in many real-world applications and competitions. However, their application in online learning scenarios, notably in reinforcement learning (RL), has been limited. In this work, we bridge this gap by introducing Gradient-Boosting RL (GBRL), a framework that extends the advantages of GBT to the RL domain. Using the GBRL framework, we implement various actor-critic algorithms and compare their performance with their NN counterparts. Inspired by shared backbones in NN we introduce a tree-sharing approach for policy and value functions with distinct learning rates, enhancing learning efficiency over millions of interactions. GBRL achieves competitive performance across a diverse array of tasks, excelling in domains with structured or categorical features. Additionally, we present a high-performance, GPU-accelerated implementation that integrates seamlessly with widely-used RL libraries (available at https://github.com/NVlabs/gbrl). GBRL expands the toolkit for RL practitioners, demonstrating the viability and promise of GBT within the RL paradigm, particularly in domains characterized by structured or categorical features.
@@ 1 Introduction
Reinforcement Learning (RL) has shown great promise in various domains that involve sequential decision making. However, many real-world tasks, such as inventory management, traffic signal optimization, network optimization, resource allocation, and robotics, are represented by structured observations with categorical or mixed data types. These tasks can benefit significantly from deployment and training on edge devices due to resource constraints. Moreover, interpretability is crucial in these applications for regulatory reasons and for trust in the decision-making process. Current neural network (NN) based solutions struggle with interpretability, handling categorical data, and supporting light implementations suitable for low-compute devices.
Gradient Boosting Trees (GBT) is a powerful ensemble method extensively used in supervised learning due to its simplicity, accuracy, interpretability, and natural handling of structured and categorical data. Frameworks such as XGBoost [7], LightGBM [20], and CatBoost [36] have become integral in applications spanning finance [49], healthcare [54, 27, 43], and competitive data science [6]. Despite their successes, GBT has seen limited application in RL. This is primarily because traditional GBT libraries are designed for static datasets with predefined labels, contrasting with the dynamic nature of RL. The distribution shift in both input (state) and output (reward) poses significant challenges for the direct application of GBT in RL. Moreover, there is a notable lack of benchmarks or environments tailored for structured data, further hindering progress in this area.
In this paper, we introduce Gradient Boosting Reinforcement Learning (GBRL), a GBT framework tailored for RL. Our contributions are:
  1. 1.
    GBT for RL. We demonstrate the viability and potential of GBT as function approximators in RL. We present GBT-based implementations of PPO, A2C, and AWR, and show that GBRL is competitive with NNs across a range of environments. In addition, similarly to supervised learning, GBRL outperforms NNs on categorical tasks (see Figure 1).
  2. 2.
    Tree-based Actor-Critic architecture. Inspired by shared architectures in NN-based actor-critic (AC), we introduce a GBT-based AC architecture. This reduces the memory and computational requirements by sharing a common ensemble structure for both the policy and value. This approach significantly reduces runtime compared to existing GBT frameworks, thus removing the barrier to solving complex, high-dimensional RL tasks with millions of interactions.
  3. 3.
    Modern GBT-based RL library  [^1] The GBRL core library is available at https://github.com/NVlabs/gbrl. The actor-critic implementations based on GBRL are available at https://github.com/NVlabs/gbrl_sb3.. We provide a CUDA-based [33] hardware-accelerated GBT framework optimized for RL. GBRL is designed to work as part of a broader system and seamlessly integrates with popular repositories such as Stable-baselines3 [39]. This new tool offers practitioners a powerful option for exploring GBT in RL settings.
Figure 1: PPO GBRL vs PPO NN. Aggregated mean and standard deviation of the normalized average reward for the final 100 episodes. Rewards were normalized as: $\text{reward}_{\text{norm}}=\frac{\text{reward}}{\text{reward}_{\text{max}\{\text{NN},\text{GBRL}\}}}$ per environment and then aggregated across each domain.
@@ 2 Related Work
Gradient boosted trees.
Recent advances have extended GBT’s capabilities beyond traditional regression and classification. In ranking problems, GBT has been used to directly optimize ranking metrics, as demonstrated by frameworks like StochasticRank [51] and recent advancements explored in Lyzhin et al. [26]. Additionally, GBT offer probabilistic predictions through frameworks like NGBoost [11], enabling uncertainty quantification [28]. The connection between GBT and Gaussian Processes [52, 45] offers further possibilities for uncertainty-aware modeling. Recently, Ivanov and Prokhorenkova [18] modeled graph-structured data by combining GBT with graph neural networks.
Despite their versatility, applying GBT in RL remains a relatively less explored area. Several works have employed GBT as a function approximator within off-policy RL methods, including its use in Q-learning [1] and in bandit settings to learn inverse propensity scores [24]. Recently, Brukhim et al. [5] proposed a boosting framework for RL where a base policy class is incrementally enhanced using linear combinations and nonlinear transformations. However, these previous works have not yet demonstrated the scalability and effectiveness in complex, high-dimensional RL environments requiring extensive interactions. In this work, we show how to adapt the framework of GBT to successfully solve large-scale RL problems.
Interpretability.
Due to the inherent non-linearities, NNs are challenging to interpret and require sophisticated methods to do so. Interpreting NNs often involves either approximation with simpler models such as decision trees or using gradient-based techniques, which require additional forward and backward passes [16, 9, 44, 3, 37]. On the other hand, interpretability methods for GBT can take advantage of the structure of a decision tree for high speed, efficiency, and accuracy [25, 10].
Structured and categorical data.
Previous work in RL has predominantly focused on using NNs due to their ability to capture complex patterns in high-dimensional data. Techniques such as Q-learning and AC methods have advanced significantly, demonstrating success in tasks involving raw sensory inputs like images, text, and audio. However, NNs that perform well on structured and categorical data typically have very specialized architectures [19, 46, 14, 2] and are not standard multi-layer perceptrons (MLPs) that are often used in many RL tasks and algorithms [34]. Even with these specialized architectures, Gradient Boosting Trees (GBT) often perform equally or better on structured and categorical datasets [19, 31, 14, 15].
@@
In this approach, the policy is parameterized by a growing linear combination of functions [29]. Each linear addition represents the functional gradient with respect to current parameters. Kersting and Driessens [21] demonstrated the direct optimization of policies using the policy gradient theorem [48]. Similarly, Scherrer and Geist [41] proposed a functional gradient ascent approach as a local policy search algorithm. While these works lay theoretical groundwork, practical results on complex, high-dimensional RL environments have not been shown. To adapt GBT’s to RL, we leverage the framework of functional gradient ascent. This combination enables a seamless integration of GBRL directly into existing RL optimization packages, such as Stable-baselines [39].
3 Preliminaries
We begin by introducing Markov Decision Processes (MDPs) and the AC schema. Then, we introduce GBT. In the following section, we show how to combine both of these paradigms into GBRL.
3.1 Markov Decision Process
@@
We consider a fully observable infinite-horizon Markov decision process (MDP) characterized by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R})$. At each step, the agent observes a state $\operatorname{\mathbf{s}}\in\mathcal{S}$ and samples an action $\operatorname{\mathbf{a}}\in\mathcal{A}$ from its policy $\pi(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$. Performing the action causes the system to transition to a new state $\operatorname{\mathbf{s}}^{\prime}$ based on the transition probabilities $P(\operatorname{\mathbf{s}}^{\prime}|\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$, and the agent receives a reward $\operatorname{\mathbf{r}}\sim\mathcal{R}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$. The objective is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted reward $J(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{\mathbf{r}}_{t}]$, with a discount factor $\gamma\in[0,1)$.
@@
The action-value function $Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}_{t}=\operatorname{\mathbf{a}}]$ estimates the expected returns of performing action $\operatorname{\mathbf{a}}$ in state $\operatorname{\mathbf{s}}$ and then acting according to $\pi$. Additionally, the value function $V_{\pi}(\operatorname{\mathbf{s}}):=\mathbb{E}_{\pi}[\sum_{t^{\prime}=0}^{\infty}\gamma^{t^{\prime}}\mathcal{R}(\operatorname{\mathbf{s}}_{t+t^{\prime}},\operatorname{\mathbf{a}}_{t+t^{\prime}})|\operatorname{\mathbf{s}}_{t}=\operatorname{\mathbf{s}}]$, predicts the expected return starting from state $\operatorname{\mathbf{s}}$ and acting according to $\pi$. Finally, the advantage function $A_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}):=Q_{\pi}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V_{\pi}(\operatorname{\mathbf{s}})$, indicates the expected relative benefit of performing action $\operatorname{\mathbf{a}}$ over acting according to $\pi$.
3.2 Actor-Critic Reinforcement Learning
Actor-critic methods are a common method to solve the objective $J(\pi)$. They learn both the policy and value. In the GBRL framework, we extend three common AC algorithms to support GBT-based function approximators.
A2C [32] is a synchronous, on-policy AC algorithm designed to improve learning stability. The critic learns a value function, $V(\operatorname{\mathbf{s}})$, used to estimate the advantage. This advantage is incorporated into the policy gradient updates, reducing variance and leading to smoother learning. The policy is updated using the following gradient: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$.
PPO [42] extends A2C by improving stability. This is achieved through constrained policy update steps using a clipped surrogate objective. This prevents drastic policy changes and leads to smoother learning. To achieve this, PPO solves the following objective: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}[\nabla_{\theta}\text{clip}(\frac{\log\pi_{\theta}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})}{\log\pi_{\theta_{\text{old}}}(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})},1-\epsilon,1+\epsilon)A(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})]$. Additionally, PPO enhances sample efficiency by performing multiple optimization steps on each collected rollout.
AWR [35] is an off-policy AC algorithm. Provided a dataset $\mathcal{D}$, AWR updates both the policy and the value through supervised learning. This dataset can be pre-defined and fixed (offline), or continually updated using the agents experience (replay buffer). At each training iteration $k$, AWR solves the following two regression problems:
$V_{k}$ $=\operatorname*{arg\,min}_{V}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\|G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})-V(\operatorname{\mathbf{s}})\|_{2}^{2}]\,,\,\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}\sim\mathcal{D}}[\log\pi(\operatorname{\mathbf{a}}|\operatorname{\mathbf{s}})\exp(\frac{1}{\beta}A_{k}(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}}))]\,,$
where $G(\operatorname{\mathbf{s}},\operatorname{\mathbf{a}})$ represents the monte-carlo estimate or TD($\lambda$) of the expected return [47].
3.3 Gradient Boosting Trees as Functional Gradient Descent
Gradient boosting trees (GBT) [12] are a non-parametric machine learning technique that combines decision tree ensembles with functional gradient descent [30]. GBT iteratively minimizes the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ over a dataset $D=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. A GBT model, $F_{K}$, predicts outputs using $K$ additive trees as follows:
$F_{K}(\mathbf{x}_{i})=F_{0}+\sum_{k=1}^{K}\epsilon h_{k}(\mathbf{x}_{i})\,,$ (1)
where $\epsilon$ is the learning rate, $F_{0}$ is the base learner, and each $h_{k}$ is an independent regression tree partitioning the feature space.
In the context of functional gradient descent, the objective is to minimize the expected loss $L(F(\mathbf{x}))=\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ with respect to the functional $F$. Here, a functional $F:\mathcal{H}\rightarrow\mathbb{R}$ maps a function space to real numbers. A GBT model can be viewed as a functional $F$ that maps a linear combination of binary decision trees to outputs: $F:\text{lin}(\mathcal{H})\rightarrow\mathbb{R}^{D}$, where $\mathcal{H}$ is the decision tree function class.
We start with an initial model, $F_{0}$, and iteratively add trees to $F$ to minimize the expected loss. Similar to parametric gradient descent, at each iteration $k$, we minimize the loss by taking a step in the direction of the functional gradient $g^{i}_{k}:=\nabla_{F_{k-1}}L(\mathbf{y}_{i},F_{k-1}(\mathbf{x}_{i}))$. However, we are constrained to gradient directions within $\mathcal{H}$. Thus, we project the gradient $g_{k}$ into a decision tree by solving:
$h_{k}=\arg\min_{h}\|-\epsilon g_{k}-h(\mathbf{x})\|_{2}^{2}\,.$ (2)

### v2

@@ Abstract
We present Gradient Boosting Reinforcement Learning (GBRL), a framework that adapts the strengths of gradient boosting trees (GBT) to reinforcement learning (RL) tasks. While neural networks (NNs) have become the de facto choice for RL, they face significant challenges with structured and categorical features and tend to generalize poorly to out-of-distribution samples. These are challenges for which GBTs have traditionally excelled in supervised learning. However, GBT’s application in RL has been limited. The design of traditional GBT libraries is optimized for static datasets with fixed labels, making them incompatible with RL’s dynamic nature, where both state distributions and reward signals evolve during training. GBRL overcomes this limitation by continuously interleaving tree construction with environment interaction. Through extensive experiments, we demonstrate that GBRL outperforms NNs in domains with structured observations and categorical features while maintaining competitive performance on standard continuous control benchmarks. Like its supervised learning counterpart, GBRL demonstrates superior robustness to out-of-distribution samples and better handles irregular state-action relationships.
Machine Learning, Reinforcement Learning, Gradient Boosting Trees, Gradient Boosting, ICML
@@ 1 Introduction
Many real-world decision-making tasks involve structured observations, where data can be organized in a tabular format and follow predefined organizational patterns. Unlike unstructured data (such as images or raw sensor data), structured observations often include heterogeneous features, both numerical and categorical. These features carry direct semantic meaning that can be used for prediction without complex feature extraction. Such structured observations closely resemble tabular datasets commonly encountered in supervised learning. Notable examples where structured observations are crucial include recommendation systems, healthcare diagnostics, digital advertising, fraud detection, weather prediction, and financial trading.
In these domains, a key challenge lies in handling structured observations, where the importance of different components varies dynamically with the agent’s state and task. Neural network (NN)-based solutions struggle with such data, requiring sophisticated architectures and preprocessing techniques to capture structural relationships (Zabërgja et al., 2024; Kadra et al., 2021; Arik & Pfister, 2021; Gorishniy et al., 2021; Hollmann et al., 2023). This often results in complex models that sacrifice sample efficiency and generalization capability.
Gradient boosting trees (GBT) offer a promising alternative through their natural handling of structured data. Their success in supervised learning stems from an iterative ensemble-building process, in which each decision tree refines predictions by partitioning the input space along individual features. This approach has made GBT frameworks such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018) integral in domains such as finance (Tian et al., 2020), healthcare (Wassan et al., 2022; Ma et al., 2022; Seto et al., 2022), and competitive data science (Chen, 2023).
This approach creates an ensemble of piecewise constant functions that excel at capturing non-smooth patterns and abrupt transitions in data (Jeffares et al., 2024; Beyazit et al., 2023; Grinsztajn et al., 2022). These properties are particularly valuable in RL, where state-action relationships often exhibit sudden changes, for instance, when transitioning between different phases of a task or when certain state variables cross critical thresholds. Such discontinuous relationships are common in structured RL tasks, from inventory management where actions change discretely with stock levels, to game scenarios where strategic decisions shift abruptly based on state conditions.
Despite these potential benefits, adapting GBT to RL presents significant challenges. Traditional GBT frameworks are designed for static datasets with stable feature distributions and predefined loss functions, where the focus is on optimizing batch training efficiency. This contrasts sharply with RL’s dynamic nature, where state distributions evolve during training, data is collected iteratively, and rewards are often delayed. While traditional GBT libraries can technically be adapted to RL using custom loss functions and incremental learning, our experiments with CatBoost and XGBoost (Section 4.3) show that this approach fails to scale due to their supervised learning-oriented design.
We address these challenges through Gradient Boosting Reinforcement Learning (GBRL), a framework that bridges the gap between GBT’s strengths and RL’s unique requirements. Our contributions are:
1. New framework for GBT in RL: We establish GBT as a viable function approximator for RL by devising popular algorithms—PPO, A2C, and AWR—with GBT backends. In popular environments, our framework demonstrates competitive performance against NNs while showing particular strengths on structured tasks.
2. Improved robustness: We demonstrate that GBT’s inherent strengths translate effectively to RL tasks. Through extensive dedicated experiments, we show superior robustness to out-of-distribution scenarios, resilience to spurious correlations, and better handling of state-space perturbations – critical advantages over NNs for real-world applications.
3. Practical implementation: We provide a CUDA-accelerated (NVIDIA, 2025) implementation that seamlessly integrates with existing RL libraries like Stable-baselines3 (Raffin et al., 2021) [^1] The GBRL core library is available at https://github.com/NVlabs/gbrl.  [^2] Actor-Critic implementations integrated within Stable-baselines3, are available at https://github.com/NVlabs/gbrl_sb3. Our shared Actor-Critic design achieves state-of-the-art performance while significantly reducing both memory and computational requirements on modern hardware.
@@ 2 Related Work
GBT Beyond Supervised Learning.
Recent advances have extended the capabilities of GBT beyond traditional regression and classification. In ranking problems, GBT has been used to directly optimize ranking metrics (Lyzhin et al., 2023), as demonstrated by frameworks such as StochasticRank (Ustimenko & Prokhorenkova, 2020). Furthermore, GBT offers probabilistic predictions through frameworks such as NGBoost (Duan et al., 2020), allowing uncertainty quantification (Malinin et al., 2021). The connection between GBT and Gaussian processes (Ustimenko et al., 2023; Sigrist, 2022) offers further possibilities for uncertainty-aware modeling. Recently, Ivanov & Prokhorenkova (2021) modeled graph-structured data by combining GBT with graph NNs.
Despite their versatility, applying GBT in RL remains a relatively less explored area. Several works have employed GBT as a function approximator within off-policy RL methods, including its use in Q-learning (Abel et al., 2016) and in bandit settings to learn inverse propensity scores (London et al., 2023). Recently, Brukhim et al. (2022) proposed a boosting framework for RL where a base policy class is incrementally enhanced using linear combinations and nonlinear transformations, forming a 2-layer NN. However, these previous works have not demonstrated scalability and effectiveness in complex, high-dimensional RL environments that require extensive interactions. In this work, we show how to adapt the framework of GBT to successfully solve large-scale RL problems for the first time.
Tabular Data.
Previous work in RL has focused predominantly on the use of NNs due to their ability to capture complex patterns in high-dimensional data. Techniques such as Q-learning and Actor-Critic methods have advanced significantly, demonstrating success in tasks involving raw sensory inputs like images, text, and audio. However, NNs that perform well on tabular data typically have very specialized architectures (Katzir et al., 2021; Somepalli et al., 2021; Gorishniy et al., 2021; Arik & Pfister, 2021) and are different from the standard multi-layer perceptrons that are often used in RL tasks and algorithms (Ota et al., 2024). Even with specialized architectures, GBT often performs equally or better on tabular datasets (Borisov et al., 2021; McElfresh et al., 2023; Grinsztajn et al., 2022; Shwartz-Ziv & Armon, 2022). GBT’s success can be attributed to a key difference in inductive biases compared to NNs. Jeffares et al. (2024) have shown that the kernel representations induced by GBTs are bounded and often behave more predictably on irregular or out-of-distribution inputs. In contrast, NN-tangent kernels can be unbounded and vary significantly on test points far from the training distribution.
@@
Kersting & Driessens (2008) introduced Non-Parametric Policy Gradients (NPPG). By combining the policy gradient theorem (Sutton et al., 1999) with functional gradient ascent, NPPG directly optimizes the policy. Although NPPG sets the foundation for combining GBT with RL, it is limited to discrete action spaces and policies represented by the Gibbs distribution. GBRL extends the ideas laid in NPPG to general Actor-Critic algorithms and generalizes to any policy representation and objective.
Figure 1: The GBRL framework. The actor’s policy and critic’s value function are parameterized by the tree ensemble $F_{k}$. For example, $F_{k}(\operatorname{s})=[\mu(\operatorname{s}),\sigma(\operatorname{s}),V(\operatorname{s})]$ for a Gaussian policy. Starting from $F_{0}$, at each training iteration, $k$, GBRL collects a rollout and computes the gradient $\nabla_{\pi_{F_{k}}}J(\pi_{F_{k}})$ with respect to the current ensemble. This gradient is then used to fit the next tree. Adding the tree to the ensemble updates it to $F_{k}(\operatorname{s})=F_{k-1}(\operatorname{s})+\epsilon h_{k}(\operatorname{s})$, where $\epsilon$ is the learning rate.
3 RL Preliminaries
@@
A fully observable infinite-horizon MDP is characterized by the tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R})$. At each step, the agent observes a state $\operatorname{s}\in\mathcal{S}$ and samples an action $\operatorname{a}\in\mathcal{A}$ from its policy $\pi(\operatorname{s},\operatorname{a})$. The action causes the system to transition to a new state $\operatorname{s}^{\prime}$ based on the transition probabilities $P(\operatorname{s}^{\prime}|\operatorname{s},\operatorname{a})$, and produce a reward $\operatorname{r}\sim\mathcal{R}(\operatorname{s},\operatorname{a})$. The agent’s objective is to maximize the expected discounted reward $J(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}]$, with a discount factor $\gamma\in[0,1)$.
@@
The state-action value function $Q_{\pi}(\operatorname{s},\operatorname{a}):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}|\operatorname{s}_{0}=\operatorname{s},\operatorname{a}_{0}=\operatorname{a}]$ estimates the expected returns of performing action $\operatorname{a}$ in state $\operatorname{s}$ and then acting according to $\pi$. Additionally, the value function $V_{\pi}(\operatorname{s}):=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}\operatorname{r}_{t}|\operatorname{s}_{0}=\operatorname{s}]$, predicts the expected return starting from state $\operatorname{s}$ and acting according to $\pi$. Finally, the advantage function $A_{\pi}(\operatorname{s},\operatorname{a}):=Q_{\pi}(\operatorname{s},\operatorname{a})-V_{\pi}(\operatorname{s})$, indicates the expected relative benefit of performing action $\operatorname{a}$ over acting according to $\pi$.
Actor-Critic methods are a common approach in RL. They simultaneously learn a policy and a value function to improve the efficiency and stability of training. For conciseness, in the body of the paper we focus on Proximal Policy Optimization (PPO) (Schulman et al., 2017) which jointly trains an actor and critic. In the supplementary material, we also showcase the implementation and results for A2C (Mnih et al., 2016) and AWR (Peng et al., 2019).